16. The town of Villageton has 500 residents, of which 300 are employed, 30 are not work- ing but actively seeking work, 80 aPlease answer a and b.
Answer #1

Given,

The the town of Villageton:

Number of employed persons = 300

Number of unemployed persons = 30 (who are not working but actively seeking work)

Note that children, retirees, full-time students and institutionalized persons are not counted under labor force (i.e., in the employed and unemployed data)

Labor force = Number of Employed + Number of unemployed = 300 + 30 = 330

a. Unemployment rate = Number of unemployed/Labor force x 100 = 30/330 x 100 = 9.09%

b. If 10 out of 30 college students graduate and enter labor force out of which only 5 find a job, then

Labor force = 330 + 10 = 340

Number of employed = 300 + 5 = 305

Number of unemployed = 30 + 5 = 35

Unemployment rate = Number of unemployed/Labor force x 100 = 35/340 x 100 = 10.29%